A large pizza at Mario’s has a 14-inch diameter and costs $12.95. If the medium
12-inch pizza costs the same per square inch, what is the cost of a medium pizza?
A. $11.10 B. $10.21 C. $9.51 D. $8.16

Answer by rfer(16322)

a=pir^2
a=3.14*7^2
a=153.86 sq in large
12.95/153.86=.084
--------------------
a=3.14*6^2
a=113.04
113.04*.084=$9.51